A new car is purchased for 19900 dollars. The value of the car depreciates at 7.5% per year. What will the value of the car be, to the nearest cent, after 8 years?


1
Expert's answer
2022-03-01T08:08:19-0500

Let P be the purchasing price =19900

Let X be the depreciated price

The rate r = 7.5%

Number of years n =8 years

X = P(1-r )n

= 19900 (1- 0.075)8

= 19900 (0.925)8

=19900 (0.53596183)

=10665.64043

The depreciated price of the car is approximately 10665.64 dollars.